OpenStudy (anonymous):

solve the equation by a method of your choice : x sq. - 2x - 3 =0 ???? if there are two solutions, separate by a comma.

OpenStudy (mertsj):

Factor it.

OpenStudy (mertsj):

\[x^2-2x-3=0\]

OpenStudy (blurbendy):

x^2 -2x - 3 = 0 (x -3)(x +1) = 0 x = 3 x = -1 When I factored, I found out what times what equals -3 but ALSO ADDS together to get -2.

OpenStudy (anonymous):

x = 3 x = -1

OpenStudy (blurbendy):

That's right!

OpenStudy (anonymous):

bye i have to go ill be right back!!!!

OpenStudy (blurbendy):

See ya'